OpenStudy (anonymous):

n/2 - 2n/16=3/8 can someone solve it.

OpenStudy (anonymous):

make it so all the fractions have the same denominator

OpenStudy (anonymous):

(n/2)*(8/8)=? (3/8)*(2/2)=?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

8/8=?

OpenStudy (anonymous):

8/8=1 you can multiply 1 to anything and still get the same number this is how you change the denominator while still keeping the fraction the same n/2 - 2n/16=3/8 will become 8n/16 - 2n/16=6/16 now multiply 16 to both sides to get rid of the denominator 8n-2n=6 solve for n